OpenStudy (anonymous):

A 10-ft diagonal brace on a bridge connects a support of the center of the bridge to a side support on the bridge. The horizontal distance that it spans is 2 ft longer that the height that it reaches on the side of the bridge. Find the horizontal and vertical distances spanned by this brace.

OpenStudy (anonymous):

This is what i have so far (x + 2)^2 + x^2 = 100 X^2 +4 + 2 (x) (2) +x^2 = 100 Using, (a + b)^2 = a^2 + b^2 +2ab 2x^2 +4x -96 = 0 So x^2 + 2x -48 = 0 Its 8 feet and 6 feet. X is 6 ft. Using this formula, Ax^2 +bx +c = 0 x = -b ± √(b^2- 4ac)/2a 8=-b±√(b^2-4ac)/2a

OpenStudy (anonymous):

from here I have no clue.

OpenStudy (anonymous):

or am i wrong.

OpenStudy (anonymous):

Lol Look at the pythagorean thereom Ill send a pic with it Solve for y [4y^2 + y^2] = (5^1/2)*y = 10 y = 10/(5)^1/2 = 4.5 ft (vertical) and 9.0ft (horizontal)

OpenStudy (anonymous):

I am sorry that is a mistake...I made the horizontal distance 2 times the vertical new solution in a minute

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

u there.